A cyclist completed 5 laps of a circular track in one hour. If his average speed was 20 miles per hour, what is the diameter of
the track

Answer by ewatrrr(24785)


Hi
A cyclist completed 5 laps of a circular track in one hour OR 1 lap in 1/5hr
Distance = r*t = 20mph*1/5hr = 4mi around the track
Circumference = pi*d
4mi = pi*d
4mi/3.14 = diameter of the track.